OpenStudy (anonymous):

what is the value of [(5+12i)^1/2 (5-12i)^1/2]

OpenStudy (anonymous):

what is this question for?

OpenStudy (anonymous):

complex no. it is a part of a question

OpenStudy (anonymous):

169 over 4

terenzreignz (terenzreignz):

hehe... \[\huge (5+12i)^{\frac12}(5-12i)^{\frac12}=[(5+12i)(5-12i)]^\frac12\] Care to do the rest, @leena1996 ?

terenzreignz (terenzreignz):

<hums> \[\huge z\cdot\bar{z}=|z|^2\]

OpenStudy (anonymous):

ya thanks for tht property....@terenzreignz

terenzreignz (terenzreignz):

I was just doodling... get back to evaluating, @leena1996 :D

terenzreignz (terenzreignz):

\[\large \sqrt i \times \sqrt i = i\quad \quad \quad ?\] Yes, @ParthKohli

terenzreignz (terenzreignz):

What?! Are you trying to confuse me on purpose?

OpenStudy (anonymous):

guys just google it you will be clear on expontial forms of iota alright dont get confused....m gonna go offline for sometime and work on more sums and then later i will come up with any questions....

OpenStudy (gorv):

@terenzreignz is correct 100 % and ans is 13

terenzreignz (terenzreignz):

Terence is awesome 100% LOL haha, can't resist Just kidding by the way :) Have a god day :D

OpenStudy (anonymous):

i havent slept in two days and completely misread omg hahaha

terenzreignz (terenzreignz):

'tis fine :D